OpenStudy (anonymous):

What is the product of the product of 2 times the square root of 35 x and 3 times the square root of 14 x? Simplify if possible

OpenStudy (anonymous):

294xsquare root of 10 42xsquare root of 10 12x2square root of 10 42square root of 10 times x squared

OpenStudy (***[isuru]***):

here's how... \[2\sqrt{35x} \times 3\sqrt{14x}\]\[(2\times \sqrt{5}\sqrt{7x}) \times (3\times \sqrt{2}\sqrt{7x})\]\[6\times7x \times \sqrt{5}\times \sqrt{2}\]\[42x \sqrt{10}\] so...... 2nd one is the answer

OpenStudy (anonymous):

thank you so much.